Determine the equation of line of distance P from the origin
1
Expert's answer
2020-01-20T09:59:20-0500

Suppose that the equation of the line is given "Ax+By+C=0" , and the distance from the origin "O(0;0)" to the straight line is P

"d=\frac{|Ax_0 +By_0+C|}{\sqrt{A^2+B^2}}, x_0=0, y_0=0, d=P\\" , then


"P=\frac{|C|}{\sqrt{A^2+B^2}},\\
C=\pm P \sqrt{A^2+B^2},"

then the equation of the line at the distance p from the origin is given by the formula

"Ax+By\pm P \sqrt{A^2+B^2}=0"

  where A, B arbitrary and "A^2+B^2 \neq 0 ."